Hi ther thanks again ur a great helppin hand .Here is the question-
the volume of a sphere is given by V=4/3TTR^3where R= the radius of the sphere. If the radius is decreased by 25%, find the change in the volume.

Answer by nerdybill(7384)

the volume of a sphere is given by V=4/3TTR^3where R= the radius of the sphere. If the radius is decreased by 25%, find the change in the volume.
.
Original volume:
V=(4/3)(pi)R^3
.
If radius = R - .25R = (3/4)R
V=(4/3)(pi)[(3/4)R]^3
V=(3/4)^3(4/3)(pi)R^3
V=(27/64)(4/3)(pi)R^3
.
Comparing the original to the one above, we find that the change in volume will be 27/64 of the original volume